# arXiv:2603.02196v1[cs.AI] 02 Mar 2026

### v1

@@ Abstract
An agent must try new behaviors to explore and improve. In high-stakes environments, an agent that violates safety constraints may cause harm and must be taken offline, curtailing any future interaction. Imitating old behavior is safe, but excessive conservatism discourages exploration. How much behavior change is too much? We show how to use any safe reference policy as a probabilistic regulator for any optimized but untested policy. Conformal calibration on data from the safe policy determines how aggressively the new policy can act, while provably enforcing the user’s declared risk tolerance. Unlike conservative optimization methods, we do not assume the user has identified the correct model class nor tuned any hyperparameters. Unlike previous conformal methods, our theory provides finite-sample guarantees even for non-monotonic bounded constraint functions. Our experiments on applications ranging from natural language question answering to biomolecular engineering show that safe exploration is not only possible from the first moment of deployment, but can also improve performance.
@@ 1 Introduction
Safe exploration is one of the most studied problems in AI safety (Amodei et al., 2016). The choice between probable safety and the hope of possible discovery presents a dilemma known by many names, from the explore-exploit tradeoff in reinforcement learning to the validity-power tradeoff in statistics. The optimal balance is problem-dependent, yet general solutions must be problem-independent. Hence to apply to specific problems any general solution must encode this balance as a hyperparameter, which then must be tuned from data or derived from assumptions about problem structure. In consequential settings, collecting data from an unsafe policy may be unacceptable. Deriving the balance from assumptions about the problem reintroduces a subtler form of problem-dependence, namely the practitioner’s understanding of the setting itself. Yet consequential, poorly understood settings are precisely where machine learning offers the greatest potential benefit.
@@
Consider an agent that observes states, takes actions, and receives both rewards and constraint feedback. The agent’s goal is to maximize reward while respecting a risk tolerance, $\alpha$. For example, a language model answering medical questions must avoid false claims, and a generative model proposing molecules must ensure they are synthesizable. A new policy has been optimized for performance but not yet tested. A safe policy already controls the risk, and instead of blindly deploying the untested policy, the agent can reuse its existing deployment data for calibration through importance weighting. The importance weights, however, depend on the deployed policy, and the deployed policy depends on the importance-weighted risk estimates. How should the agent proceed? Rather than assuming knowledge of the problem, we resolve this circularity by exploiting what the agent already knows about itself.
@@
We introduce conformal policy control (CPC, Figure 1), a method for safe exploration that uses the calibration data to interpolate between any safe and any optimized policy while provably respecting the risk threshold. CPC requires no assumptions on the reward or constraint functions, no access to the optimized policy’s training process, and no additional samples to tune the balance beyond what a safe policy already provides. The key idea is to parameterize the balance as a likelihood-ratio threshold between the safe and optimized policies. Conformal calibration on the safe policy’s data determines the most aggressive threshold that still respects the declared risk. At deployment, the interpolated policy is realized through rejection sampling, which allows the agent to probabilistically self-regulate to a zone of competence determined by its calibration data. Because CPC operates entirely at test time, the same safe and optimized policies can be reused under different risk tolerances, trading test-time compute for risk control without retraining.
@@
Providing finite-sample guarantees in this setting poses several challenges for uncertainty quantification. The deployed policy depends on the calibration data, introducing subtle, data-dependent distribution shift. The relevant losses are general constraint functions, not just coverage indicators. Conformal risk control (CRC) (Angelopoulos et al., 2024) addresses both of these challenges, but assumes a monotone relationship between the control parameter and the loss. In CPC, the constraint function does not depend on the control parameter, so the monotonicity assumption does not hold. We extend CRC to non-monotonic bounded loss functions, and we prove that CPC inherits finite-sample guarantees when the control parameter governs the policy, not the loss.
@@
We validate our method on three tasks. In medical question answering, we control the false discovery rate, a non-monotonic loss that standard CRC cannot handle directly. In constrained active learning, the agent’s actions induce feedback-loop shifts that violate standard exchangeability assumptions. In black-box sequence optimization, a language model agent must improve biomolecular sequences while respecting a constraint budget. In each setting, CPC provides finite-sample risk control from the first round of deployment, without hyperparameter tuning or structural assumptions. In medical question answering, this yields tighter risk control and better claim recall than existing methods. In active learning, we obtain finite-sample guarantees where previous approaches provide only asymptotic control. In sequence optimization, we extend conformal methods to a setting where no prior guarantees existed, and find that moderate risk control can counterintuitively improve optimization performance by stabilizing the algorithm. Safe exploration is not only safe, it can also be more efficient.
2 Background
@@
We consider a decision-making agent that interacts with an environment over a sequence of rounds. At each round $t$, the agent observes a context (i.e., prompt) $X_{t}\in\mathcal{X}$ and selects an action $A_{t}\in\mathcal{A}$ according to a policy $\pi(A\mid X)$. If the policy is parameterized by an LLM, then $\mathcal{A}\subseteq\mathcal{X}$. The agent then observes (potentially noisy) rewards $R_{t}\in\mathbb{R}$ and losses $L_{t}\in\mathbb{R}$. The reward function $r(x,a)$ quantifies the degree to which a context-action pair is aligned with the agent’s objective, and the loss function $\ell(x,a)$ quantifies whether the context-action pair satisfies the agent’s constraints. While a policy can be defined implicitly by solving an optimization problem with respect to estimated reward and loss surrogates (Watkins and Dayan, 1992; Jones et al., 1998), in this paper we focus on controlling explicit policies parameterized by a conditional distribution (Williams, 1992), such as an autoregressive transformer.
@@ 2.1 The Safe Policy Improvement Problem
Post-training design decisions immediately prompt two questions: which policy divergence measure is best, and how much divergence is allowable to maintain an acceptable risk profile? The first question has been studied extensively (see Section 3). The second, however, is typically treated as a hyperparameter tuning problem (Bergstra et al., 2011; Feurer and Hutter, 2019), in which the user must specify a constraint on a “control parameter”—such as a divergence budget or penalty weight—and the algorithm optimizes subject to that constraint. The resulting indirection creates a gap between what the user actually wants and the control parameters that are accessible in available algorithms (Hutchins et al., 1986; Norman, 1988). The user’s goal is declarative (i.e., a desired outcome of the program): control risk at some level $\alpha$. The algorithms demand imperative instructions (i.e., the steps the program should execute): bound divergence at some level, or set the penalty weight to some value. Without a principled translation, the user must learn the mapping from outcome to instruction by trial and error, spending the very samples the agent seeks to use efficiently.
@@ 2.2 Conformal Risk Control
Ideally, the user would specify a risk threshold $\alpha$ and the algorithm would automatically determine the corresponding policy constraint. Conformal risk control (CRC; Angelopoulos et al., 2024) provides exactly this capability, but for a restricted class of problems. While CRC does not directly solve the policy control problem described above, it provides foundational ideas we build upon.
CRC is a framework for setting a control parameter in a way that controls the expected value of a user-specified loss function with finite-sample guarantees. The key assumptions are that the calibration loss functions $L_{1},\ldots,L_{n}$ and test loss $L_{n+1}$ are exchangeable, and that each loss function $L(\lambda)$ is monotonically nonincreasing with the control parameter $\lambda$. CRC selects $\hat{\lambda}$ to be the smallest value such that the empirical risk on the calibration data falls below the target level adjusted by a finite-sample correction, which enables the guarantee $\mathbb{E}[L_{n+1}(\hat{\lambda})]\leq\alpha$. Conformal prediction (Vovk et al., 2005) is recovered as a special case when the loss is the miscoverage indicator $L(\lambda)=\mathbbm{1}\{Y\notin\hat{C}_{\lambda}(X)\}$ for labels $Y$ and prediction sets $\hat{C}_{\lambda}$, where the control parameter $\lambda$ controls the size of the prediction set. When the calibration and test distributions differ, weighted conformal methods (Tibshirani et al., 2019) can be used to restore validity by reweighting the empirical risk according to the likelihood ratio $p_{\text{test}}/p_{\text{cal}}$.
The primary limitation of CRC for policy optimization is that the loss function must be monotonically nonincreasing with a tunable control parameter $\lambda$. Many losses of practical interest, such as indicators of violating feasibility/safety sets, cannot be parameterized in this way—unlike in the case of conformal prediction described above, where one can set the size of the prediction set to modulate the loss value, feasibility/safety sets are typically fixed. Our approach addresses this limitation by introducing the control parameter in the policy or data distribution that the expectation is taken over, rather than in the loss function. Even with this modification, the resulting expected loss is not necessarily monotonic in the control parameter, which presents a theoretical challenge we discuss further in Section 4.
@@ 3 Abbreviated Related Work
Conservative model-based optimization: A standard approach to safe policy improvement relies on the observation that the risk of a new policy can be bounded in terms of the risk of a reference policy plus a penalty that grows with the divergence between them (Kakade and Langford, 2002). When the reference policy is known to satisfy the safety constraint, controlling this divergence provides a mechanism for controlling risk. The idea of bounding risk through bounded divergence motivates a broad family of locality-constrained policy optimization methods (Kim et al., 2025). Prior work has proposed entropy-regularized control and KL-penalized objectives (Todorov, 2009; Fox et al., 2016), trust-region policy optimization (Schulman et al., 2015, 2017), conservative value penalties in offline reinforcement learning (Kumar et al., 2020; Trabucco et al., 2021), trust-region methods in black-box optimization (Eriksson et al., 2019), and uncertainty-penalized safe Bayesian optimization (Sui et al., 2015). Damani et al. (2023) proposed using a likelihood ratio as a safety filter for model-based optimization, but treated the likelihood ratio threshold as a hyperparameter for empirical tuning. [^2] Bounding the likelihood ratio $\pi_{t}(a\mid x)/\pi_{0}(a\mid x)\leq\beta$ uniformly implies a bound on the KL divergence and other $f$-divergences between $\pi_{t}$ and $\pi_{0}$. In this sense, likelihood ratio clipping is a strict form of divergence constraint.
From an optimization perspective, our work is most similar to Stanton et al. (2023), who proposed constraining a Bayesian optimization policy to regions where the resulting conformal prediction sets would have bounded width given a declarative miscoverage tolerance $\alpha$. However, their approach was limited by three weaknesses. First, their approach was restricted to controlling the miscoverage loss. Second, their use of an implicit policy created an awkward fixed-point problem: finding the policy that induced an action distribution under which the coverage would be controlled at the desired level. Third, they did not fully address the multi-step feedback covariate shift induced by sequential model-based optimization. [^3] A form of covariate shift in which the test distribution is dependent on previously observed data. Indeed, Stanton et al. (2023) noted that, at the time, the existing conformal theory had not yet proved general enough guarantees to cover the multi-round model-based optimization setting.
@@
However, the techniques used throughout this line of work do not enable one to prescribe a data distribution—e.g., select a policy—that guarantees coverage. That is, these methods can provide prediction sets with coverage guarantees for actions generated by any fixed policy (a “descriptive” use of conformal techniques), but using those prediction sets to then select a policy (a “prescriptive” use) introduces further, unaccounted-for dependence between the observed data and test policy that nullifies coverage guarantees.
To enable this prescriptive use case, we build upon techniques from conformal risk control (CRC) (Angelopoulos et al., 2024), a general framework for selecting a hyperparameter with finite-sample guarantees for any monotonic loss, of which conformal prediction can be cast as a special case. When the loss of interest is non-monotonic, Angelopoulos et al. (2024) also showed that deploying the procedure on a monotonic relaxation of the losses provides asymptotic guarantees of control. However, they left the problem of finite-sample guarantees for non-monotonic losses open. In the next section, we develop a variant of CRC that drops the monotonicity requirement, and instead relies on smoothness and a form of stability of the losses to control the risk on the test point. Finally, we apply techniques from Prinster et al. (2024) to prove finite-sample guarantees for our procedure in the multi-round model-based optimization setting. See Appendix A for an extended discussion of related work.
Concurrently with this paper, Angelopoulos (2026) also presents CRC guarantees for non-monotonic losses. A primary difference is that paper studies parameterized losses, whereas our paper introduces and focuses on policy control, where the control parameter tunes the policy rather than the loss. Even in the comparable setting of parameterized losses, however, the contributions are distinct, with neither subsuming the other: Relative to the univariate method in Angelopoulos (2026), our method (Section 4.1) will always be at least as a-priori conservative (e.g., as in Figure 3, right); consequently, whereas that paper’s analogous guarantee relies on leave-one-out stability, our results use replace-one stability, which is generally a more relaxed assumption (Bousquet and Elisseeff, 2002; Shalev-Shwartz et al., 2010).

### v2

# Conformal Policy Control

@@ Abstract
An agent must try new behaviors to explore and improve. In high-stakes environments, an agent that violates safety constraints may cause harm and must be taken offline, curtailing any future interaction. Imitating old behavior is safe, but excessive conservatism discourages exploration. How much behavior change is too much? We show how to use any safe reference policy as a probabilistic regulator for any optimized but untested policy. Conformal calibration on data from the safe policy determines how aggressively the new policy can act, while provably enforcing the user’s declared risk tolerance. Unlike conservative optimization methods, we do not assume the user has identified the correct model class nor tuned any hyperparameters. Unlike previous conformal methods, our theory provides finite-sample guarantees even for non-monotonic bounded loss functions. Our experiments on applications ranging from natural language question answering to biomolecular engineering show that safe exploration is not only possible from the first moment of deployment, but can also improve performance.
@@ 1 Introduction
Safe exploration is one of the most studied problems in AI safety (Amodei et al., 2016). The choice between probable safety and the hope of possible discovery presents a dilemma known by many names, from the explore-exploit tradeoff in reinforcement learning to the validity-power tradeoff in statistics. The optimal balance is problem-dependent, yet general solutions must be problem-independent. So, to apply to real-world problems, any general solution must encode this balance as a hyperparameter, which then must be tuned from data or derived from assumptions about problem structure. In consequential settings, collecting data from an unsafe behavior policy (a context-dependent strategy or distribution over actions) may be unacceptable. Deriving the balance from assumptions about the problem reintroduces a subtler form of problem-dependence, namely the practitioner’s understanding of the setting itself. Yet consequential, poorly understood settings are precisely where machine learning offers the greatest potential benefit.
@@
Consider an agent that observes contexts (i.e., states) and takes actions to maximize reward while satisfying safe operating constraints to within some tolerance. For example, a language model responding to medical questions aims to give answers that are as helpful as possible while avoiding false claims. In biomolecular engineering, a generative model proposing improvements to lead molecules aims to maximize efficacy of its designs while ensuring that they are synthesizable. In such settings, the agent has candidate policies optimized for performance but does not yet know which, if any, are safe to deploy. From past experience, the agent usually has a default safe policy it knows controls the risk, but playing it safe may be far from optimal.
What the safe policy does supply is calibration data: Reweighting each observation by a candidate-to-safe likelihood ratio gives an estimate of each candidate’s risk. But which candidate should the agent deploy? The natural move is to estimate every candidate’s risk and pick the most aggressively optimized one that still looks safe, but this approach introduces bias: Selection corrupts the procedure by favoring estimates that understate their true risk. Yet the change in risk the agent must control is due to the policy shift it will author. The agent does not know how the world will respond to a given action, but it knows exactly how each candidate would distribute its actions, which is the only difference between the distributions it has experienced in the past and those it would encounter in the future. [^1] If we assume the environment is stationary.
@@
We introduce conformal policy control (CPC, Figure 1), a method for safe exploration that uses the calibration data to interpolate between any safe and any optimized policy while provably respecting the risk tolerance. CPC requires no assumptions on the reward or constraint functions, no access to the optimized policy’s training process, and no additional samples to tune the balance beyond what a safe policy already provides. The key idea is to invert the usual importance-weighting paradigm: Rather than estimating risk for a fixed policy, we search over candidate policies, parameterized by a likelihood-ratio threshold, for the most aggressive one such that all evaluated candidates still satisfy the risk tolerance (with a weighted conservative adjustment). Conformal calibration on the safe policy’s data determines both the weights and the threshold. At deployment, the interpolated policy can be realized through rejection sampling (or other Monte Carlo methods), which allows the agent to probabilistically self-regulate to a zone of competence determined by its calibration data. Because CPC operates entirely at test time, the same safe and optimized policies can be reused under different risk tolerances, trading test-time compute for risk control without retraining.
@@
Providing finite-sample guarantees in this setting poses several challenges for uncertainty quantification. The deployed policy depends on the calibration data, introducing subtle, data-dependent distribution shift. The relevant losses are general constraint functions, not just coverage indicators. Conformal risk control (CRC) (Angelopoulos et al., 2024) can address these challenges, but assumes a monotone relationship between the control parameter and the loss. In CPC, the loss function does not depend on the control parameter, so the monotonicity assumption does not hold. We generalize CRC (gCRC) to non-monotonic bounded loss functions, and we prove that CPC inherits finite-sample guarantees when the control parameter governs the policy, not the loss.
@@
We validate our methods on three tasks. In medical question answering, gCRC controls the false discovery rate, a non-monotonic loss that standard CRC cannot handle, achieving tighter control and better claim recall than existing methods. In constrained active learning, CPC provides finite-sample guarantees under feedback-loop covariate shifts where previous approaches offered only asymptotic control. In black-box biomolecular sequence optimization, CPC extends conformal methods to a setting where no prior guarantees existed, and we find that moderate risk control can counterintuitively improve optimization performance by reducing wasted evaluations on infeasible actions. Safe exploration is not only safe, it can also be more efficient.
2 Preliminaries and Background
@@
Notation: We consider a decision-making agent that interacts with an environment over a sequence of rounds. At each round $t$, the agent observes a context (i.e., prompt) $X_{t}\in\mathcal{X}$ and selects an action $A_{t}\in\mathcal{A}$ according to a policy $\pi(A\mid X)$. If the policy is parameterized by an LLM, then $\mathcal{A}\subseteq\mathcal{X}$. The agent then observes (potentially noisy) rewards $R_{t}\in\mathbb{R}$ and losses $L_{t}\in\mathbb{R}$. Throughout, uppercase letters denote random variables and lowercase letters their realized values. The reward function $r(x,a)$ quantifies the degree to which a context-action pair is aligned with the agent’s objective, and the loss function $\ell(x,a)$ quantifies the degree to which the context-action pair violates the agent’s constraints. While a policy can be defined implicitly by solving an optimization problem with respect to estimated reward and loss surrogates (Watkins and Dayan, 1992; Jones et al., 1998), in this paper we focus on controlling explicit policies parameterized by a conditional distribution (Williams, 1992), such as an autoregressive transformer.
@@ 2.1 The Safe Policy Improvement Problem
Post-training design decisions immediately prompt two questions: which policy divergence measure is best, and how much divergence is allowable to maintain an acceptable risk profile? The first question has been studied extensively (see Section 3). The second, however, is typically treated as a hyperparameter tuning problem (Bergstra et al., 2011; Feurer and Hutter, 2019), in which the user must specify a constraint on a “control parameter” (such as a divergence budget or penalty weight) and the algorithm optimizes subject to that constraint. The resulting indirection creates a gap between what the user actually wants and the control parameters that are accessible in available algorithms (Hutchins et al., 1986; Norman, 1988). The user’s goal is declarative (i.e., a desired outcome of the program): control risk at some level $\alpha$. The algorithms demand imperative instructions (i.e., the steps the program should execute): bound divergence at some level, or set the penalty weight to some value. Without a principled translation, the user must learn the mapping from outcome to instruction by trial and error, spending the very samples the agent seeks to use efficiently.
As previewed in the introduction, importance weighting offers a more direct route. Given data from the safe policy, the agent can estimate the expected loss under any candidate policy by reweighting each observation by the likelihood ratio between the candidate and safe policies, via the identity
$\operatorname*{\mathbb{E}}_{\begin{subarray}{c}X\sim p(\cdot)\\ A\sim\pi_{t}(\cdot\mid X)\end{subarray}}[\ell(X,A)]=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}X\sim p(\cdot)\\ A\sim\pi_{0}(\cdot\mid X)\end{subarray}}\left[\frac{\pi_{t}(A\mid X)}{\pi_{0}(A\mid X)}\,\ell(X,A)\right].$
Bounding this ratio at some threshold $\beta$ simultaneously defines a constrained policy and bounds the importance weights, so the safe policy’s data suffices to determine whether the candidate still respects the risk tolerance. [^3] Bounding the likelihood ratio $\pi_{t}(a\mid x)/\pi_{0}(a\mid x)\leq\beta$ uniformly implies a bound on the KL divergence and other $f$-divergences between $\pi_{t}$ and $\pi_{0}$. In this sense, likelihood ratio clipping is a strict form of divergence constraint.
@@ 2.2 Conformal Risk Control
What remains is a principled way to choose $\beta$ from finite safe-policy data, with a guarantee on the risk of the resulting deployed policy. Conformal risk control (CRC; Angelopoulos et al., 2024) provides exactly this for a restricted class of problems: it sets a control parameter $\lambda$ to control the expected value of a user-specified loss $L_{i}(\lambda)$. Given exchangeable (e.g., IID) calibration loss functions, $L_{1}(\cdot),\ldots,L_{n}(\cdot)$, and test loss function, $L_{n+1}(\cdot)$, the key assumption is that each $L_{i}(\lambda)$ is monotonically nonincreasing with the control parameter, $\lambda$. For example, conformal prediction (Vovk et al., 2005) is a special case of CRC where the loss is the miscoverage indicator $L(\lambda)=\mathbbm{1}\{Y\notin\hat{C}_{\lambda}(X)\}$ for labels $Y$ and prediction sets $\hat{C}_{\lambda}$, and $\lambda$ controls the size of the prediction set: increasing the prediction set size can only decrease miscoverage. CRC extends conformal prediction to cases such as setting an inclusion threshold on predicted probabilities for multilabel classification with false negative rate (FNR) control (i.e., controlling the expected proportion of true labels that are not included in the set). Like miscoverage, FNR is monotonic in $\lambda$: adding labels to the prediction set can only increase the proportion of true labels contained. By selecting $\hat{\lambda}$ as the smallest value such that the (conservatively adjusted) empirical risk falls below $\alpha$, CRC attains the finite-sample guarantee $\mathbb{E}[L_{n+1}(\hat{\lambda})]\leq\alpha$. In each case, increasing $\lambda$ can only decrease the loss. This monotonicity is what lets CRC search from aggressive to conservative and stop at the first value that satisfies the guarantee.
For policy control, however, it is not immediately clear how CRC could be applied. Many losses of practical interest admit no tunable control parameter at all. Consider again the feasibility indicator $L_{t}:=\mathbbm{1}\{A_{t}\notin\mathcal{F}\}$: a prediction set $\hat{C}_{\lambda}$ can be grown to reduce miscoverage, but the feasibility set $\mathcal{F}$ is fixed by the environment, and there is no $\lambda$ that makes a given action more or less feasible. We instead place the control parameter on the policy, governing the distribution of actions rather than the loss they incur. Even so, the resulting expected loss is not necessarily monotonic in that parameter, a theoretical challenge we address in Section 4.
@@ 3 Abbreviated Related Work
Conservative model-based optimization: A standard approach to safe policy improvement relies on the observation that the risk of a new policy can be bounded in terms of the risk of a reference policy plus a penalty that grows with the divergence between them (Kakade and Langford, 2002). When the reference policy is known to satisfy the safety constraint, controlling this divergence provides a mechanism for controlling risk. The idea of bounding risk through bounded divergence motivates a broad family of locality-constrained policy optimization methods (Kim et al., 2025). Prior work has proposed entropy-regularized control and KL-penalized objectives (Todorov, 2009; Fox et al., 2016), trust-region policy optimization (Schulman et al., 2015, 2017), conservative value penalties in offline reinforcement learning (Kumar et al., 2020; Trabucco et al., 2021), trust-region methods in black-box optimization (Eriksson et al., 2019), and uncertainty-penalized safe Bayesian optimization (Sui et al., 2015). Damani et al. (2023) proposed using a likelihood ratio as a safety filter for model-based optimization, but treated the ratio threshold as a hyperparameter.
From an optimization perspective, our work is most similar to Stanton et al. (2023), who proposed constraining a Bayesian optimization policy to regions where the resulting conformal prediction sets would have bounded width given a declared miscoverage tolerance $\alpha$. However, their approach was limited by three weaknesses. First, their approach was restricted to the miscoverage loss. Second, their use of an implicit policy created an awkward fixed-point problem: finding the policy that induced an action distribution that controlled the coverage. Third, they did not fully address the multi-step feedback covariate shift induced by sequential model-based optimization. [^4] A form of covariate shift in which the test distribution is dependent on previously observed data. Indeed, Stanton et al. (2023) noted that, at the time, the existing conformal theory had not yet proved general enough guarantees for that setting.
@@
However, the techniques used throughout this line of work do not enable one to prescribe a data distribution (e.g., select a policy) that guarantees coverage. That is, these methods can provide prediction sets with coverage guarantees for actions generated by any fixed policy (a “descriptive” use of conformal techniques), but using those prediction sets to then select a policy (a “prescriptive” use) introduces further, unaccounted-for dependence between the observed data and test policy that nullifies coverage guarantees.
To enable this prescriptive use case, we build upon techniques from conformal risk control (CRC) (Angelopoulos et al., 2024), a general framework for selecting a hyperparameter with finite-sample guarantees for any monotonic loss, of which conformal prediction can be cast as a special case. When the loss of interest is non-monotonic, Angelopoulos et al. (2024) also showed that deploying the procedure on a monotonic relaxation of the losses provides asymptotic guarantees of control. However, they left the problem of finite-sample guarantees for non-monotonic losses open. In the next section, we develop a variant of CRC that drops the monotonicity requirement, and instead relies on smoothness of the losses and a form of stability of the calibration procedure to control the risk on the test point. Finally, we apply techniques from Prinster et al. (2024) to prove finite-sample guarantees for our procedure in the multi-round model-based optimization setting. See Appendix A for an extended discussion of related work.
Figure 4: Rejection sampling for policy interpolation at three $\beta$ values. As $\beta$ increases from near-zero to large values, the constrained policy $\pi_{t}^{(\beta)}$ interpolates from the safe policy $\pi_{0}$ to the optimized policy $\pi_{t}$. Blue histogram bars show accepted samples matching $\pi_{t}^{(\beta)}$, while teal/red bars show rejected samples from the proposal distribution. (a) Small $\beta$ nearly recovers the safe policy (b) Intermediate $\beta$ shows intermediate interpolation with frequent rejections. (c) Large $\beta$ approaches the optimized policy with minimal rejection.
Concurrently, Angelopoulos (2026) also present CRC guarantees for non-monotonic losses, but for parameterized losses, rather than the policy-control setting we introduce here where the control parameter tunes the policy instead of the loss. Even in the comparable setting of parameterized losses, however, the contributions are distinct, with neither subsuming the other: Relative to the univariate method in Angelopoulos (2026), our method (Section 4.1) will always be at least as conservative a priori (e.g., as in Figure 3, right); consequently, whereas that paper’s analogous guarantee relies on leave-one-out stability, our results use replace-one stability, which is generally a more relaxed assumption (Bousquet and Elisseeff, 2002; Shalev-Shwartz et al., 2010).